Two cars leave from the same place at the same time and drive in opposite directions. One car travels an average of 6 mph faster than the other. After 2 hours, the cars are 310 miles apart. Find the speed of each car.
Answer by checkley77(12569)

2X+2(X+6)=310
2X+2X+12=310
4X=310-12
4X=298
X=298/4
X=74.5 mph for the slower car.
74.5+6=80.5 mph for the faster car.
proof:
2*74.5+2*80.5=310
149+161=310
310=310